[figure: 9.7 cm]

EVOLUTION FROM BCS TO BOSE LIMIT IN A D-WAVE SUPERCONDUCTOR

L.S. Borkowski

Institute of Physics, A. Mickiewicz University, Umultowska 85, 61-614 Poznań, Poland

We study the evolution from BCS to Bose limit in a two-dimensional d-wave superconductor at zero temperature. We examine single particle properties when particle density and attraction strength are varied. A dramatic change of these properties occurs when the system approaches the bosonic limit.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : L.S. Borkowski

Address for correspondence : Instytut Fizyki UAM Umultowska 85 61-614 Poznań

Email address : lsb@amu.edu.pl